# Asymmetric Influence of the Amplitude-Dependent Tune Shift on the Transverse Mode-Coupling Instability

### v1

@@ IV.5 Betatron Tune Shift with Current
Overall, it can be concluded that, for the presented measurements, the sign of the ADTS leads to an asymmetry in how strongly the vertical TMCI is self-containing and in which tune shift with current is observed above the threshold. As expected, a higher absolute value of the ADTS coefficient leads to a lower maximal center-of-mass oscillation and bunch size blow up. But when comparing signs, a higher positive ADTS coefficient is required for the same amount of suppression of the instability than for a negative ADTS.
@@
Compared to previous studies on using Landau damping together with ADTS to mitigate the TMCI, two differences were found for the presented investigations for the parameters at the 3 GeV ring at MAX IV. It is clear that, the Landau damping comes in as a stabilization mechanism only after the beam begins to go unstable, as the threshold was shown to not be dependent on the ADTS coefficient, leading to the observed saw-tooth pattern. The second result and rather interesting finding is that the maximal level that the center-of-mass motion and bunch size reaches before being contained by Landau damping is higher for a positive than a negative sign of the ADTS coefficient. This could not only be shown in the measurement but also in tracking simulations and in theoretical stability considerations including Landau damping and the ADTS as well.
@@
A key contribution to making this asymmetry visible, might by the rather low synchrotron frequency at the 3 GeV ring due to the low momentum compaction factor, typical in fourth-generation light-source storage rings combined with the low RF frequency of 100 MHz. Both of these aspects lead to a synchrotron frequency that starts to be within the betatron tune spread of the bunch when it is blown up by the instability. This would indicate that the presented findings might become more relevant in new fourth-generation light-source storage rings with even more extreme parameters than MAX IV.

### v2

@@ IV.5 Betatron Tune Shift with Current
Our experiments and analysis show an asymmetry of how strongly the vertical TMCI is self-containing and of which tune shift with current is observed above the threshold. As expected, a higher absolute value of the ADTS coefficient leads to a lower maximal center-of-mass oscillation and bunch size blow-up. But when comparing signs, a higher positive ADTS coefficient is required for the same amount of suppression of the instability than for a negative ADTS.
@@
Compared to previous studies on using Landau damping together with ADTS to mitigate the TMCI, two differences were found for the presented investigations for the parameters at the 3 GeV ring at MAX IV. It is clear that, the Landau damping only comes in as a stabilization mechanism after the beam has become unstable, as the threshold was shown to not be dependent on the ADTS coefficient, resulting in the observed saw-tooth pattern.
The second result and rather interesting finding is that the maximal level that the center-of-mass motion and bunch size reaches before being contained by Landau damping is higher for a positive than a negative sign of the ADTS coefficient. This could not only be shown in the measurement but also in tracking simulations and in theoretical stability considerations including Landau damping and the ADTS as well. The intuitive explanation behind this is that the ADTS contributes asymmetrically to the Landau damping depending on the ADTS sign, as the resulting tune spread either covers the same lower frequencies as the tune shift with current by impedance or it goes in the other direction towards higher frequencies reducing the resulting Landau damping effect significantly as seen in the presented results for positive ADTS coefficients.
@@
A key contribution for this asymmetry to become relevant and in making it observable, might be the rather low synchrotron frequency at the 3 GeV ring. The low momentum compaction factor, typical for fourth-generation light-source storage rings, is combined with the low RF frequency of 100 MHz. Both of these aspects lead to a synchrotron frequency that then lies within the betatron tune spread of the bunch when it is blown up by the instability. This therefore puts the betatron tune at the point of mode-coupling within the reach of the Landau damping before the beam hits the vacuum chamber wall and is lost. The role of the low synchrotron frequency in enabling these dynamics indicates that the presented findings will become more relevant in future fourth-generation light-source storage rings with even more extreme parameters than MAX IV.